OpenStudy (asad786):

will medal Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (asad786):

i think b

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-(-a)=+a

OpenStudy (anonymous):

z-(-0.3)=-2.4 z+0.3=-2.4 z=-2.4-0.3=-2.7